OpenStudy (anna):

what is the value of theta here,, 4sin2theta=0

OpenStudy (sid1729):

\[4\times \sin (2\theta) = 0\] If you divide both sides by 4, you get: \[\frac{4 \times \sin(2\theta)}{4} = \frac{0}{4}\] which reduces to \[\sin (2\theta) = 0\] Now, what are the values where the sin function can be zero? These values are : 0,\[\pi, 2\pi, 3\pi, 4\pi.....npi\] where n is a natural number Thus, \[2\theta = 0, \pi, 2\pi, 3\pi, 4\pi, .... npi\] You can divide both sides by two to get the answer.